Rosa is 14 years older than Amy. Three years ago, she was 3 times as old as Amy. How old is each girl now?
Answer by rm29924(97)

Let Amy's current age = X
Rosa's current age = X + 14
Amy's age three years ago = X - 3
Rosa's age three years ago = 3(X-3)
Rosa's current age minus three should equal Rosa's age three years ago
X + 14 - 3=3(X-3)
X+11=3X-9
x+20=3x
20=2x
x=10

Amy's current age is 10 years old
Rosa's current age is 24 years old
check answer:
10+14-3=3(10-3)
24-3=30-9
21=21